OpenStudy (anonymous):

simplify 5 1/3 + (-3 9/18)

OpenStudy (whpalmer4):

Dr. @Hero to the OR, please :-)

OpenStudy (whpalmer4):

We did one yesterday, see here: http://openstudy.com/study#/updates/50f0c97ee4b0abb3d86f6c71

OpenStudy (jkristia):

you can change 5 1/3 to somthing/6 and you can change -3 9/18 to something/6, then you subtract the fractions

OpenStudy (anonymous):

i do not know

OpenStudy (jkristia):

change 1/3 to ?/6 => (2*1)/(2*3) => 2/6. Now 6/6 is the same as 1, so if you have 5 (ones) you have (5*6)/6 = 30/6, so 5 1/3 = 32/6, now do the same for -3 9/18

OpenStudy (anonymous):

hmmmmm

OpenStudy (anonymous):

11/6 or 1 5/6 is the answer